Balance Sheet Accruals Cash& MS Receivables Inventories Fixed Assets 30,000 50,000 Accou Common Equity Annual Sales$1,500,000 Net Income $60,000 Cost of Goods Sold-$1,200,000 Number of Shares Outstanding-20,000 OC What is the P/E multiple? O 9.5 O 11.1 15.0 18.0 20.0 QUESTION 48
Answer #1

Earnings per share = Net income / number of shares outstanding

Earnings per share = 60,000 / 20,000

Earnings per share = 3

P/E = Price / EPS

P/E = 60 / 3

P/E = 20.0